OpenStudy (anonymous):

The Scale Of A Map Is 1 in. : 10 mi. How Many Actual Miles Does Each Measurement On The Map Represent.? 1. 3 1/4 2. 2 1/2 ^(Make Into Decimal)^ 3. 4.4 4. 5.3

OpenStudy (anonymous):

So 2. Would Be 2.5

OpenStudy (anonymous):

3. 3.25

OpenStudy (texaschic101):

1. 32.5 miles 2. 25 miles 3. 44 miles 4. 53 miles

OpenStudy (texaschic101):

Let me show you how I did these... 1/10 = 3.25/x (1 in to 10 mi = 3.25 in to x miles) cross multiply (1)(x) = (3.25)(10) x = 32.5 miles That is the first one...you do the same with the others

OpenStudy (anonymous):

Yes, That's What I Got Too.(: I Just Multiplied Them By 10. @texaschic101

OpenStudy (texaschic101):

Basically that was all you had to do. Just multiply them by 10. Good job :)

OpenStudy (anonymous):

Thank You Tho.(: I Started Working The Problems Out Then Realized I Had This Open And You Had Done The Same Thing I Did. c: @texaschic101

OpenStudy (texaschic101):

Great minds think alike...lol :)